708 - (257 - (125 - 841) x 283)= ?
1
Expert's answer
2012-10-01T08:24:54-0400
708 - (257 - (125 - 841) x 283)=708 - 257 + (125 - 841) x 283=708 - 257 + (125 - 841) x 283=451-716*283=-202177